OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (−1, 6) and has a slope of −3. A.y − 6 = −3x − 3 B.y − 6 = −3(x + 1) C.y = −3x + 3 D.3x + y = 3

OpenStudy (anonymous):

@ali2x2 @DanJS

OpenStudy (anonymous):

can i get some help plz

OpenStudy (anonymous):

can someone help me

OpenStudy (anonymous):

can u help me @mathway and @dan815

OpenStudy (anonymous):

Do you still need help?

OpenStudy (anonymous):

Use the formula. \[\large\ y-y _{1}=m(x-x_{1})\] where m= -3 y_1=6 y_2=-1